Question 2: Should you reduce the price or increase advertising?

The selling price is $20/unit, variable costs are $10/unit, and fixed costs are $3,000 in total. Sales volume decreased to 200 units because of a recession.
You are considering two options to stimulate sales:
(1) Reduce the price to $18/unit. This will increase sales volume by 20%.
(2) Buy additional advertising for $300 and keep the original price. This will increase sales volume by 20%.

Use the gross approach to decide whether you should do nothing (the status quo), reduce the price, or increase advertising.

status quo (1) reduce the price (2) increase advertising
Volume in units
Revenue $ $ $
   Variable costs $ $ $
Contribution margin $ $ $
  Fixed costs $ $ $
Profit* $ $ $

* enter losses as a negative number: e.g., a loss of $500 should be entered as -500, not as (500) or ($500).

What should you do?

Reduce the price

Do nothing   

Increase advertising
Answer #1

Answer:

(1) Reduce in Selling Price

Selling Price is $18 per unit

Sales volume = 200 * 120% = 240 units

Variable cost is $10 per unit, Total variable cost is $2,400 (240 * $10)

Fixed cost is $3,000

(2) Increase in advertising

Selling Price is $20 per unit

Sales volume = 200 * 120% = 240 units

Variable cost is $10 per unit, Total variable cost is $2,400 (240 * $10)

The adverting expense is fixed cost. So,fixed cost is $3,300

Status quo (1) reduce the price (2) Increase advertising
A Volume in Units 200 240 240
B Selling Price per Unit $20 $18 $20
C Revenue (A * B) $4,000 $4.320 $4,800
D Less: Variable Costs $2,000 $2,400 $2,400
E Contribution Margin (C-D) $2,000 $1,920 $2,400
F Less: Fixed Cost $3,000 $3,000 $3,300
G Profit /Loss (E-F) -1,000 -1,080 -900

By considering the Gross approach, it better to Increase the advertising.

Because, even if we reduce the price or do nothing the loss we are going to incur is more compared to Increase advertising

So, it is better to increase advertising as it is the most viable option among the three alternatives.